OpenStudy (anonymous):

A tank, with a rectangular base and having vertical sides of height 4 m, is initially full of water. The water leaks out of a small hole in the horizontal base of the tank at a rate which, at any instant, is proportional to the square root of the depth of the water at that instant. If x is the depth of water at time t after the leak started, write down a differential equation connecting x and t. If the tank is exactly half empty after one hour, find the further time that elapses before the tank becomes completely empty

OpenStudy (phi):

"water leaks out" means change in volume. Volume of the tank is the area of the base times its height. Call the area of the base B. Its height is 4 m. So V= 4B x is the depth of water at time t. So the volume of water in the tank at time t is V= Bx The change in volume wrt to time is \[\frac{dV}{dt}= B\frac{dx}{dt}\] "leak is proportional to the square root of the depth" k is some unknown constant. \[ B\frac{dx}{dt}= k x^{1/2}\] separate the variables: \[ x^{-1/2} dx = \frac{k}{B} dt \]